The resistance of 120 meters of wire is 16Ω. What is the conductance of the wire from one end to the middle?
1
Expert's answer
2021-09-20T19:10:13-0400

Gives

l=120m

R=16 ohm

We know that

"\rho=\frac{RA}{l}\\\sigma=\frac{l}{RA}"

Put value

"\sigma=\frac{120}{16\times A}=\frac{7.5}{A} mho/m"